OpenStudy (anonymous):

QUESTION 7 OF 10 A single die is rolled. Find the probability of rolling an even number or a number less than 6.

OpenStudy (anonymous):

2/9

OpenStudy (kropot72):

The probability of rolling an even number is found by noting that there are three even numbers 2, 4 and 6 each with a probability of being rolled equal to 1/6. therefore the combined probability of rolling an even number is: \[\frac{1}{6}+\frac{1}{6}+\frac{1}{6}=\frac{3}{6}\] Thye numbers less than 6 are 1, 2, 3, 4 and 5. But 2 and 4 have already been considered in calculating the probability of rolling an even number. The probability of rolling a 1, 3 or 5 is 3/6 as found for the 3 even numbers. The combined probabilities are therefore: \[\frac{3}{6}+\frac{3}{6}=1\]